OpenStudy (anonymous):

If f(x) = x 2 + 1, find f(a + 1).

OpenStudy (anonymous):

possible answers: a 2 + 2 a 2+ a + 1 a 2+ 2a + 2

OpenStudy (anonymous):

f(x) = (x^2)/(x + 1) = (x^2 - 1 + 1)/(x + 1) = (x - 1) + 1/(x + 1) f '(x) = 1 - 1/(x + 1)² f ''(x) = 2/(x + 1)^3using quotient rule, first derivative = x(x+2)/(x+1)^2 second derivative = 2/(x+1)^3 plug 3 in you get 2/64 or 1/32F"(x)=2 /(x+1)^3 so f"(3)=2/64=1/32

OpenStudy (anonymous):

some twos are squares

OpenStudy (triciaal):

f(x) =x^2 + 1 find F(a +1) let x = a + 1 (a +1)^2 + 1 a^2 +2a + 1 + 1 a^2 + 2a + 2 the 3rd choice